# Towards Understanding the Robustness of Diffusion-Based Purification: A Stochastic Perspective

### v2

@@ 6.2 Defense Performance Under Different Conditions
Effectiveness of ADDT on different DBP models. We apply ADDT to a set of diffusion models and apply DiffPure-style DBP with the refined models. The comparison on clean and robust accuracy with the baseline and other DBP models is presented in Section 6.1. It shows that ADDT effectively enhances the robustness of these models.
@@
Model Training samples Clean $l_{\infty}$ $l_{2}$
DDPM None 85.94 47.27 69.34
@@
Performance on different classifiers. We evaluate the cross-model protection ability of ADDT fine-tuned models by applying the diffusion model trained with WRN-28-10 guidance to other classifiers. The results in Section 6.1 indicate that the adversarial purification ability of these diffusion models could be transferred to different classifiers with various architectures. Notably, using a DPEDM with WRN-28-10 Guidance training, we achieve $69.63\%$ $l_{\infty}$ robust accuracy on a WRN-70-16 classifier. This demonstrates the feasibility of ADDT as it does not require classifier-specific fine-tuning.
Performance under acceleration. Speeding up the diffusion process by omitting intermediate steps has become a common practice in the use of diffusion models (Nichol & Dhariwal, 2021; Song et al., 2020a). Hence, we evaluate the robustness of accelerated DBP models. The computation cost is measured by the number of neural function evaluations (NFE), which indicates the number of evaluation steps performed during the DBP backtracking process. For our experiments, we set $t^{*}=0.1\times T$ and accelerate the process by excluding intermediate time steps. For example, with an NFE of 5, the time steps for the DBP reverse process would be $t=[100,80,60,40,20,0]$. The results in Section 6.1 validate the effectiveness of ADDT in improving the robustness of accelerated DPDDPM models. Note that the performance of DPDDPM varies significantly between different values of NFE. This may be explained by the fact that DDPM introduces stochasticity (Gaussian noise) at each reverse step; with fewer reverse steps, its stochasticity reduces. Additionally, the generation capability of DDPM is sensitive to skipping of intermediate steps. We also conducted an evaluation of DPDDIM models, as detailed in Appendix H.
@@ 6.3 Ablation Study and Analysis
RBGM. We compare the generative ability of diffusion models fine-tuned with RBGM-mapped
Table 6: FID score of DDPM for CIFAR-10 fine-tuned to different perturbations (the lower the better). Fine-tuning with RBGM-mapped perturbations yields lower FID scores than $l_{\infty}$ perturbations (without RBGM).
@@
and $l_{\infty}$ perturbations by comparing their Fréchet Inception Distance (FID) scores (Heusel et al., 2017), as shown in Table 6. The results show that diffusion models fine-tuned with RBGM-mapped perturbations maintain generation quality comparable to the vanilla diffusion model, while models directly fine-tuned with $l_{\infty}$ perturbations without RBGM show degraded performance. We also observe that training with RBGM-mapped perturbations generalized better to different attacks. Experimental details and additional tests are presented in Appendix K.
$White$$DW$0010101010202020203030303040404040505050506060606070707070808080809090909010010010010047.2747.2747.2747.2716.8016.8016.8016.8042.1942.1942.1942.194.984.984.984.98without ADDTAccuracy (%)DPDDPMDPDDIM
$White$$DW$0010101010202020203030303040404040505050506060606070707070808080809090909010010010010051.4651.4651.4651.4639.1639.1639.1639.1646.4846.4846.4846.4817.0917.0917.0917.09with ADDT
Figure 6: Revisiting robustness under Deterministic White-box setting. ADDT improves robustness under both White-box and Deterministic White-box setting, implying that ADDT strengthens the models’ ability to purify adversarial inputs.
CGPO. We analyze the effect of fine-tuning using different training samples in Section 6.2. Specifically, we compare the performance of samples generated with classifier guidance in the CGPO step, referred to as “CGPO”, against those generated with Mean Squared Error (MSE) loss, noted as “MSE-guided”. The evaluation results are presented for DDPM with NFE 100 and DDIM with NFE 10. Results demonstrates that samples generated by CGPO significantly outperform MSE-guided samples in enhancing DBP robustness.
Revisiting DBP robustness. We re-examine robustness under the Deterministic White-box setting by comparing the performance of diffusion models with and without ADDT fine-tuning, as shown in Figure 6. The fine-tuned models show significantly higher robust accuracy under the DW-box setting, indicating improved non-stochasticity-based robustness brought by ADDT. We provide a more detailed experiment considering different NFE in Appendix M. We also compare the loss landscapes of ADDT fine-tuned models and vanilla diffusion models, as shown in Figure 3. This comparison shows that our method effectively smooths the loss landscape of DBP models and enhance its ability to purify adversarial perturbations.
Table 7: Robust accuracy (%) on CIFAR-10 under more PGD and EoT iterations.
Model PGD200+EoT20 PGD20+EoT10
@@
Evaluation with stronger PGD+EoT attacks. To balance computational cost and attack strength, we primarily employ the PGD20+EoT10 configuration in our evaluations. To further validate the efficacy of ADDT under stronger attack settings, we assess its performance using the more challenging PGD200+EoT20 setup. The results presented in Table 7 and Table 9 show that under these intensified attacks, ADDT’s robust accuracy experiences a moderate 5% drop compared to the PGD20+EoT10 setting. Nonetheless, across various settings and datasets, ADDT consistently demonstrates superior robust accuracy to the baseline.
@@ 6.4 Scaling to More Complex and High-Dimensional Data
To evaluate the scalability of DBP and ADDT on more complex and high-dimensional datasets, we extend our experiments to include Tiny-ImageNet (Le & Yang, 2015) and ImageNet-1k (Deng et al., 2009). For Tiny-ImageNet, we trained the diffusion model from scratch for 200 epochs, followed by fine-tuning with ADDT for an additional 50 epochs, guided by a pretrained WRN-28-10 classifier. For ImageNet-1k, the diffusion model was trained from scratch for 12 epochs and then fine-tuned with ADDT for 8 epochs, using a pretrained ResNet-101 classifier as guidance.
As shown in Table 9 and Table 9, ADDT successfully enhances the robustness of DBP on these complex datasets, while the improvement may be limited. Similar to the characteristics of adversarial training on classifiers, effective up-scaling of ADDT may require sufficient model capacity and a large amount of data, and our results in Section 6.1 have demonstrated the benefits of applying a larger diffusion model in DBP. However, efficient training methods specialized for DBP models can be a promising direction for future studies.
In addition, we observe that the strong EoT attacks on images of higher resolution are computationally intensive. Specifically, our evaluation with PGD200+EoT20 on 1024 images of the size $224\times 224$ requires approximately 7 days on 8 NVIDIA RTX 4090 GPUs. Therefore, we argue that the up-scaling in data dimension can also imply significantly increased computational costs for the attacker.
@@
Model Vanilla ADDT
Clean $l_{\infty}$ $l_{2}$ Clean $l_{\infty}$ $l_{2}$
@@
Table 9: Clean and robust accuracy (%) on ImageNet-1k with ResNet-101 classifier. All experiments are conducted under $l_{\infty}$ perturbation bound of $\epsilon=4/255$.
@@ 6.5 Discussions on Improving Stochasticity-based DBP Robustness
As analyzed in Section 4.2, the DBP robustness can be primarily attributed to the high variance of the stochastic attack gradients. We argue that increasing the variance of attack gradients can improve the stochasticity-based robustness of DBP models by reducing the effectiveness of EoT attacks. Specifically, on the one hand, higher variance means higher errors in the estimation of the expected attack gradient direction with a fixed number of samples, and to reduce the error, more EoT steps are required. On the other hand, higher variance also suggests that the expected deviation of the DW-box attack gradient (which suggests the most effective attack direction) deviates more from the EoT attack gradient, even if the estimation of the mean attack gradient is accurate. As discussed in Section 4.2, such deviation leads to lower increase in classification loss for one attack step, suggesting that a successful attack may not be achieved or require more PGD steps.
To increase the variance of attack gradients, an intuitive approach is to introduce more stochasticity. As an initial experiment, we augment the DBP framework’s stochasticity by integrating a Corrector sampler. Specifically, Song et al. (2021) develop a Predictor-Corrector (PC) sampler framework. While standard VPSDE (DDPM++) implementations typically use only the predictor component, we add a Corrector sampler to increase stochasticity in the reverse diffusion process, thereby boosting the overall variance of attack gradients. As detailed in Appendix J, our preliminary results indicate that this modification improves the robustness of DBP models against adaptive White-box attacks. However, there is a trade-off: the model’s clean accuracy decreases slightly. These observations align with the findings of Nie et al. (2022), where randomizing the diffusion timesteps also leads to robustness improvements at the cost of clean accuracy, as well as with prior research on stochastic preprocessing defenses (Gao et al., 2022).
@@ 7 Conclusion
This study offers a new perspective on the robustness of Diffusion-Based Purification (DBP), emphasizing the crucial role of stochasticity and challenging the traditional view that robustness is mainly derived from minimizing the distribution gap through the forward diffusion process. We introduce a Deterministic white-box (DW-box) attack scenario and show that DBP models are based on stochastic elements to evade effective attack directions and lack the ability to purify adversarial perturbations, demonstrating distinct properties compared to models trained with Adversarial Training. To further enhance the robustness of DBP models, we develop Adversarial Denoising Diffusion Training (ADDT) and Rank-Based Gaussian Mapping (RBGM). ADDT integrates adversarial perturbations into the training process, while RBGM trims perturbations to more closely resemble Gaussian distributions. Experiments across various diffusion methods, attack settings, and datasets suggest the effectiveness of ADDT. In summary, this study highlights the decoupling of stochasticity-based and purification-based robustness of DBP models for deeper analysis, and suggests combining them for better robustness in practice.
@@ Appendix A Influence of EoT Iterations on DBP Robustness Evaluation
See Figure 7 for a visual explanation, where the red line shows the DBP accuracy after attack, and the blue line shows the similarity between the attack directions of the White-box-EoT and Deterministic White-box. The trend is clear: more EoT iterations lead to greater similarity and lower model accuracy, the rate of increase in similarity and the rate of decrease in accuracy both tend to slow down with further iterations.
Balancing computational cost and evaluation accuracy, we chose the PGD20-EoT10 configuration for our robustness evaluation.
@@
Appendix B DBP Models Employing Different Stochastic Elements Cannot Be Attacked All at Once
Previous research has questioned whether stochasticity can improve robustness, arguing that it can produce obfuscated gradients that give a false sense of security (Athalye et al., 2018). To investigate this, we implement DWSemi-box, a semi-stochastic setting that restricts the stochastic elements to a limited set of options. Our results show that stochasticity can indeed improve robustness, even when the attacker has full knowledge of all the possible options for stochastic elements.
Building on the concept of Deterministic White-box, we further propose DWsemi-128 to explore whether stochasticity can indeed improve robustness. Unlike under Deterministic White-box, where the attacker attacks a DBP model under the exact set of stochastic noise used in the evaluation, DWsemi-128 relaxes the stochastic elements to a limited set of options, the attacker should simultaneously attack over 128 different sets of stochastic noise. It uses the average adversarial direction from these 128 noise settings (EoT-128) to perturb the DBP model. To understand the impact of stochasticity, we analyze the changes of the model loss under DW-box attack and DWsemi-128 attack. We plot these changes by adjusting a factor $k$ to modify an image $x$ with a perturbation $\sigma$, evaluating the loss at $x+k\sigma$ where $k$ varies from $-16$ to $16$. We generate perturbations with $l_{\infty}$ Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015) with magnitude $1/255$. The plot is evaluated using WideResNet-28-10 with DPDDPM over the first 128 images of CIFAR-10 dataset.
@@ Appendix B DBP Models Employing Different Stochastic Elements Cannot Be Attacked All at Once
Figure 8: Impact of stochasticity on perturbation efficacy. Perturbations created under DWsemi-box setting are less potent compared to DW-box setting. For non-adversarial perturbations, we randomly assign each element a value of either $1/255$ or $-1/255$.
As Figure 8 shows, in the Deterministic White-box setting, the perturbations significantly increase the loss, proving their effectiveness. However, for DWsemi-128, where the attack spans multiple noise setting, the increase in loss is more moderate. This suggests that even when the attackers are fully informed about the stochastic noise choices, stochasticity still improves the robustness of the DBP. This challenges the notion that there exists a vulnerable direction that is effective for all stochastic noise.

### v3

@@ 6.2 Defense Performance Under Different Conditions
Performance across different DBP models. We evaluate the effectiveness of ADDT by applying it to various diffusion models and comparing their performance with existing methods. The results shown in Section 6.1 indicate that ADDT consistently enhances model robustness while maintaining competitive accuracy on clean data.
@@
DBP method Training samples $Clean$ $\ell_{\infty}$ $\ell_{2}$
DPDDPM None 85.94 47.27 69.34
@@
Performance across different downstream classifiers. We evaluate ADDT’s cross-classifier protection performance by applying a diffusion model specifically ADDT fine-tuned with WRN-28-10 guidance, to protect a variety of classifiers. The results in Section 6.1 show that these models effectively protect classifiers with diverse architectures. Notably, employing a DPEDM model trained with WRN-28-10 guidance achieves a $69.63\%$ $\ell_{\infty}$ robust accuracy on a WRN-70-16 classifier. This ability to protect different classifiers without classifier-specific fine-tuning highlights the effectiveness and feasibility of ADDT.
Performance under acceleration. Speeding up the diffusion process by omitting intermediate steps has become a common practice in diffusion models (Song et al., 2020; Nichol & Dhariwal, 2021). In this context, we evaluate the robustness of accelerated DBP models. To quantify computational cost, we utilize the number of Neural Function Evaluations (NFEs), which reflects the number of evaluation steps executed during the DBP reverse process. For our experiments, we set $t^{*}=0.1\times T$ and accelerate the process by excluding intermediate time steps. For example, with 5 NFEs, the time steps for the DBP reverse process would be $t=[100,80,60,40,20,0]$. The results in Section 6.1 validate the effectiveness of ADDT in improving the robustness of accelerated DPDDPM models. Note that the performance of DPDDPM varies significantly with different NFE values. This may be because DDPM introduces stochasticity (Gaussian noise) at each reverse step, and with fewer reverse steps, the influence of this stochasticity diminishes. Additionally, the generative performance of DDPM is sensitive to the omission of intermediate steps. We also evaluated DPDDIM models, as detailed in Appendix H.
@@ 6.3 Ablation Study and Analysis
RBGM. We compare the generative ability of diffusion models fine-tuned from the same pre-trained
Table 6: FID score of DDPM fine-tuned on CIFAR-10 with different perturbations (lower is better). Fine-tuning with RBGM-mapped perturbations results in lower FID scores compared to $\ell_{\infty}$ perturbations (without RBGM).
@@
models, using two different perturbations: RBGM-mapped perturbations and $\ell_{\infty}$ perturbations. To quantify generation quality, we use Fréchet Inception Distance (FID) scores (Heusel et al., 2017), as shown in Table 6. The results demonstrate that diffusion models fine-tuned with RBGM-mapped perturbations maintain generation quality comparable to the vanilla diffusion model, while models directly fine-tuned with $\ell_{\infty}$ perturbations without RBGM show degraded performance. Additionally, we observe that training with RBGM-mapped perturbations generalizes better to different types of attacks. Experimental details and further tests are provided in Appendix L.
$White$$DW$005050505010010010010047.2747.2747.2747.2716.8016.8016.8016.8042.1942.1942.1942.194.984.984.984.98without ADDTAccuracy (%)DPDDPMDPDDIM
$White$$DW$005050505010010010010051.4651.4651.4651.4639.1639.1639.1639.1646.4846.4846.4846.4817.0917.0917.0917.09with ADDT
Figure 6: Revisiting robustness under Deterministic White-box setting. ADDT improves robustness under both White-box and Deterministic White-box settings, implying that ADDT strengthens the models’ ability to purify adversarial inputs.
CGPO. We analyze the impact of fine-tuning with different training samples in Section 6.2. Specifically, we compare the performance of samples generated with classifier guidance in the CGPO step, referred to as “CGPO”, against those generated with Mean Squared Error (MSE) loss, denoted as “MSE-guided”. The evaluation results are presented for DDPM with 100 NFEs and DDIM with 10 NFEs. The results demonstrate that samples generated by CGPO significantly outperform MSE-guided samples in terms of enhancing DBP robustness.
Revisiting DBP robustness. We re-examine robustness under the Deterministic White-box setting by comparing the performance of diffusion models with and without ADDT fine-tuning, as shown in Figure 6. The fine-tuned models show significantly higher robust accuracy under the DW-box setting, indicating that ADDT improves non-stochasticity-based robustness. Further experiments across different models and NFEs, presented in Appendix M, confirm these improvements in robustness. We also compare the loss landscapes of ADDT fine-tuned models and vanilla diffusion models, as shown in Figure 3. This comparison reveals that our method effectively smooths the loss landscape of DBP models, enhancing their ability to resist adversarial perturbations.
Evaluation with stronger PGD+EoT attacks. To balance computational cost and attack strength, we primarily employ the PGD20+EoT10 configuration in our evaluations. To further validate the efficacy of ADDT under stronger attack settings, we assess its performance using the more challenging PGD200+EoT20 setup. The results presented in Table 7 and Table 9 show that under these intensified attacks, ADDT’s robust accuracy experiences a moderate 5% drop compared to the PGD20+EoT10 setting. Nevertheless, across various configurations and datasets, ADDT consistently outperforms the baseline in terms of robust accuracy.
Table 7: Robust accuracy (%) on CIFAR-10 under stronger attacks.
DBP method PGD200+EoT20 PGD20+EoT10
@@ 6.4 Scaling to More Complex and High-Dimensional Data
To assess the scalability of DBP and ADDT on more complex datasets, we extend our experiments to include Tiny-ImageNet (Le & Yang, 2015) and ImageNet-1k (Deng et al., 2009). For Tiny-ImageNet, we trained the diffusion model from scratch for 200 epochs and then fine-tuned it with ADDT for 50 epochs, using a pretrained WRN-28-10 classifier for guidance. For ImageNet-1k, the diffusion model was trained from scratch for 12 epochs, followed by 8 epochs of ADDT fine-tuning, with guidance from a pretrained ResNet-101 classifier.
As shown in Table 9 and Table 9, ADDT improves the robustness of DBP on these complex datasets. However, the additional robustness provided by ADDT appears relatively modest. This suggests that, to scale ADDT more effectively, adequate model capacity and a sufficient volume of data are essential. This is further supported by the results in Section 6.1, where the larger DiffPure+ADDT model demonstrates greater robustness compared to the smaller DPDDPM+ADDT model, aligning with findings from traditional AT (Wang et al., 2023; Huang et al., 2023).
It is also important to note that scaling to high-resolution images poses a significant challenge for attackers, as performing strong EoT attacks on high-resolution images is computationally expensive. For instance, our evaluation with PGD200+EoT20 on $224\times 224$ images at a resolution of 1024 takes approximately 7 days of computation on 8 NVIDIA RTX 4090 GPUs.
@@
DBP method Vanilla ADDT
$Clean$ $\ell_{\infty}$ $\ell_{2}$ $Clean$ $\ell_{\infty}$ $\ell_{2}$
@@
Table 9: Clean and robust accuracy (%) on ImageNet-1k with ResNet-101 classifier. Experiments are conducted under $\ell_{\infty}$ perturbation bound of $\epsilon=4/255$.
@@ 6.5 Discussions on Improving Stochasticity-based DBP Robustness
As discussed in Section 4.2, DBP robustness can be primarily attributed to the high variance of the stochastic attack gradients. We argue that increasing the variance of attack gradients can improve the stochasticity-based robustness of DBP models by reducing the effectiveness of EoT attacks. Specifically, on one hand, higher variance leads to larger errors in estimating the expected attack gradient direction. To reduce these errors, more EoT steps are required. On the other hand, higher variance also means that the DW-box attack gradient (which indicates the most effective attack direction) deviates more from the EoT attack gradient, even if the estimation of the mean attack gradient is accurate. As discussed in Section 4.2, this deviation leads to a lower increase in classification loss after one attack step, suggesting that a successful attack may not be achieved or require more PGD steps.
To increase the variance of attack gradients, a natural approach is to introduce more stochasticity. In an initial experiment, we augment the DBP framework’s stochasticity by integrating a Corrector sampler. Specifically, Song et al. (2021b) propose a Predictor-Corrector (PC) sampler framework. While standard VPSDE (DDPM++) implementations typically use only the predictor component, we add a Corrector sampler to increase stochasticity in the reverse diffusion process, thereby enhancing the overall variance of attack gradients. As detailed in Appendix J, our preliminary results indicate that this modification improves the robustness of DBP models against adaptive White-box attacks. However, there is a trade-off: the model’s clean accuracy decreases slightly. These observations align with the findings of Nie et al. (2022), where randomizing the diffusion timesteps also improves robustness at the expense of clean accuracy, as well as with previous research on stochastic preprocessing defenses (Gao et al., 2022).
@@ 7 Conclusion
This study provides a new perspective on the robustness of Diffusion-Based Purification (DBP), highlighting the critical role of stochasticity and challenging the traditional view that robustness primarily arises from minimizing the distribution gap through the forward diffusion process. We introduce a Deterministic white-box (DW-box) attack setting and demonstrate that DBP models rely on stochastic elements to evade effective attack directions but lack the ability to purify adversarial perturbations. This shows distinct differences compared to models trained with Adversarial Training. To further improve the robustness of DBP models, we propose Adversarial Denoising Diffusion Training (ADDT) and Rank-Based Gaussian Mapping (RBGM). ADDT integrates adversarial perturbations into the training process, while RBGM trims perturbations to better approximate Gaussian distributions. Experiments across various diffusion methods, attack settings, and datasets confirm the effectiveness of ADDT in enhancing robustness. In conclusion, this study emphasizes the decoupling of stochasticity-based and purification-based robustness of DBP models for deeper analysis, and suggests that combining both approaches can lead to improved robustness in practice.
Acknowledgements
This work is supported in part by National Natural Science Foundation of China (NSFC) under Grant No.62376292, U21A20470, Guangdong Provincial General Fund No. 2024A1515010208.
We would like to express our sincere gratitude to Hongjun Wang, Zhaoyu Chen, Huanran Chen, Ruoxin Chen and Conghan Yue for their invaluable discussions, which greatly contributed to the development of this work.
@@ Appendix A Influence of EoT Iterations on DBP Robustness Evaluation
See Figure 7 for a visual explanation, where the red line shows the classification accuracy, and the blue line shows the similarity between the attack directions of the White-box-EoT and Deterministic White-box. The results show a clear trend: increasing the EoT iterations raises the similarity between the attack directions and reduces model accuracy.
Note that both the increase in similarity and the decline in accuracy converge as the number of iterations increases. Balancing computational cost and evaluation accuracy, we adopted the PGD20-EoT10 configuration for our robustness evaluation.
@@
Appendix B DBP Models with Different Stochastic Elements Cannot Be Attacked Simultaneously
Previous research has raised concerns about whether stochasticity enhances robustness, suggesting it may create obfuscated gradients that provide a false sense of security (Athalye et al., 2018a). To investigate this, we implement DWSemi-box, a semi-stochastic framework that restricts the available stochastic elements to a limited set of options, in order to explore whether stochasticity can genuinely improve robustness.
DWsemi-128 builds on the concept of Deterministic White-box. In contrast to the Deterministic White-box approach, where the attacker exploits the exact stochastic noise used during evaluation, DWsemi-128 limits the stochastic elements to a constrained set of possibilities. To mount an attack across 128 distinct sets of stochastic noise, the attacker can employ the average adversarial direction across these 128 noise settings (EoT-128) to perturb the DBP model. We evaluate the impact of stochasticity by comparing the loss changes under DW-box and DWsemi-128 attacks. This is done by adjusting a factor $k$ to modify an image $\bm{x}$ with a perturbation $\bm{\sigma}$, and evaluating the loss at $\bm{x}+k\bm{\sigma}$ for $k$ ranging from $-16$ to $16$. Perturbations are generated using the $\ell_{\infty}$ Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015) with a magnitude of $1/255$. The experiment is conducted using the WideResNet-28-10 with DPDDPM over the first 128 images from the CIFAR-10 dataset.
@@ Appendix B DBP Models with Different Stochastic Elements Cannot Be Attacked Simultaneously
Figure 8: Impact of stochasticity on perturbation efficacy. Perturbations created under the DWsemi-box setting are less effective compared to those under the DW-box setting. For non-adversarial perturbations, each element is randomly assigned a value of either $1/255$ or $-1/255$.
As shown in Figure 8, in the Deterministic White-box setting, perturbations lead to a significant increase in loss, demonstrating their effectiveness. However, under DWsemi-128, the loss increase is more moderate. This suggests that even when stochastic elements are limited to a small set of possibilities, and the attacker has fully considered all options, stochasticity still contributes to the robustness of the DBP model. These results contradict the hypothesis that there might be a universally vulnerable direction when DBP applies different stochastic elements, thereby challenging concerns about potential insecurity.